OpenStudy (anonymous):

Help! The weight of 100 identical samples of a substance is 0.01 pound. What is the weight of 1000 samples?

OpenStudy (anonymous):

0.01/100 = 0.0001 x 1000 = 0.1lb

OpenStudy (anonymous):

huhhh??

OpenStudy (anonymous):

if 100 samples ---> 0.01 lbs then 1000 samples ---> x? cross multiply and solve for x ^_^

OpenStudy (anonymous):

ok..hmm

OpenStudy (anonymous):

100x = 1000(0.01) x = 1000(0.01)/100 ^_^

OpenStudy (anonymous):

did you get the way?

OpenStudy (anonymous):

do i multiply 0.01 & 10?

OpenStudy (anonymous):

100x=0.01. Therefore, x=0.01/100=0.0001 so x * 1000 = 0.1lb

OpenStudy (anonymous):

0.01 x 1000 then divide by 100 ^_^

OpenStudy (anonymous):

0.1?

OpenStudy (anonymous):

excellent!

OpenStudy (anonymous):

pounds right?

OpenStudy (anonymous):

nvm. got it haha .h/o i got another one

OpenStudy (anonymous):

wait it said that was wrong.

OpenStudy (anonymous):

it did? are you sure?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

The weight of identical samples of a substance is 1 pound. What is the weight of samples?

OpenStudy (anonymous):

I'm sure of my answer

OpenStudy (anonymous):

wait... The weight of 1000 identical samples of a substance is 1 pound. What is the weight of 10 samples?

OpenStudy (anonymous):

answer 0.1 pounds???

OpenStudy (anonymous):

I'm sure

OpenStudy (anonymous):

its 0.01 ..

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

1000 x 0.01 = 10 10/100 = 0.1

OpenStudy (anonymous):

this is what it says...now this is what i'm dealing with. it explains it sooo confusing...u ready?lol

OpenStudy (anonymous):

it says the weight of 1000 samples is one pound. to find the weight of 1 sample we divide 1 by 1000. ---> weight of 1 sample = 1 divided by 1000= 0.0001 The weight of 10 samples is 10 times this amount. weight of 10 samples = 10xweight of one sample ___> 10x0.0001=0.01 pound.

OpenStudy (anonymous):

but you wrote that it's 0.01 lb

OpenStudy (anonymous):

intially in the question lol

OpenStudy (anonymous):

the question was The weight of 1000 identical samples of a substance is 1 pound. What is the weight of 10 samples?

OpenStudy (anonymous):

LOL! see what you wrote in the question >_< you said 0.01 !

OpenStudy (anonymous):

ohhhhh my bad! wow.

OpenStudy (anonymous):

sorry!

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

hahahah

OpenStudy (anonymous):

it;s okay :)

OpenStudy (anonymous):

the volume of 1000 drops of a liquid is 10 fl. oz. What is the volume of 10 drops?

OpenStudy (anonymous):

and im sure thats the question this time LOL

OpenStudy (anonymous):

do it the same way I solved the previous question: if 1000 drops --->10fl 10 drops -----> x? cross multiply and solve for x ^_^ try

OpenStudy (anonymous):

like 1000/10 X 10/x ?

OpenStudy (anonymous):

lol no, cross multiply and you'll get: 1000x = 10(10) x = 100/1000 x = 0.1 ^_^

OpenStudy (anonymous):

draw an X between them and you'll get 10(10) = (1000)x

OpenStudy (anonymous):

^_^ did you get it?

OpenStudy (anonymous):

yeah but im not sure why u divide 100/1000

OpenStudy (anonymous):

to solve for x, you have : 1000x = 100 right? now get x alone and we do that by dividing both sides by 1000 x = 100/1000 :)

OpenStudy (anonymous):

is it clearer now? :)

OpenStudy (anonymous):

got it!!!

OpenStudy (anonymous):

great :)

OpenStudy (anonymous):

The weight of 1000 identical samples of a substance is 0.01 pound. What is the weight of 10,000 samples? ok i had this... 1000x/ 1000 = 10/1000 x= 0.01 and it said it was wrong.